Convert 1 newton into dyne using method of dimension

Grade:12

2 Answers

Yuvraj Singh
2 years ago
Convert 1 newton into dyne using method of dimension

Ans:
the dimensional formula of ForceF=[M1L1T−2]
Soa=1,b=1,c=−2
Now because Newton is in m.k.s system, we have
M1​=1kg
L1​=1m
T1​=1s
M2​=1g
L2​=1cm
T2​=1s
n1​=1Newton
n2​=?
Using the formula
n2​=n1​[M1/M2]a[L1/L2]b[T1/T2]c
n2​=1[1kg/1g]1[1m/1cm]1[1s/1s]−2

n2​=1[1000g/1g]1[100cm/1cm]1×1

n2​=1000×100=100000=105dynes
Yuvraj Singh
2 years ago
the dimensional formula of ForceF=[M1L1T−2]
Soa=1,b=1,c=−2
Now because Newton is in m.k.s system, we have
M1​=1kg
L1​=1m
T1​=1s
M2​=1g
L2​=1cm
T2​=1s
n1​=1Newton
n2​=?
Using the formula
n2​=n1​[M1/M2]a[L1/L2]b[T1/T2]c
n2​=1[1kg/1g]1[1m/1cm]1[1s/1s]−2

n2​=1[1000g/1g]1[100cm/1cm]1×1

n2​=1000×100=100000=105dynes